Two flasks of equal volume connected by narrow tube(of negligible volume) are at 300K and contains 0.7 mole of H2(.35 mole in each flask) at 0.5 atm.One of the flask is then immersed in a bath kept at 400k while the other remains at 300K. Caculate the final pressure and the number of moles of H2 in each flask.

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
Let the two flasks be “a” & “b”. Also, let flask “a” be immersed in a bath at temperature,Ta = 127°C = 127 + 273.15 = 400.15 Kand flask “b” remains the same as initial temperature i.e., temperature,Tb = 27°C + 273.15 = 300.15 K. No. of mole for flask a and b be “na” & “nb” respectively.

Also, the final pressure in both the flask Pa and Pb will be same i.e.,Pa = Pb = Pand alsoVa = Vb = V.

UsingIdeal Gas Lawfor the combined system initially, we have

PV = nRT

⇒ 0.5 * 2V = 0.7 * R * 300.15 …[both the flasks have the same volume]

V = 210.105 R….. (i)

For flask a :

Pa * Va = na * R * Ta

⇒ P * V = na * R * Ta

⇒ P * 210.105 R = na * R * 400.15 ….. [ value of V from (i)]

na = 0.525P….. (ii)

For flask b:

Pb * Vb = nb * R * Tb

⇒ nb = (Pb Vb) / (R Tb)

⇒ nb = (P * 210.105 R) / (R * 300.15 )

nb = 0.7P….. (iii)

We know,

Total no.of moles initially was = na + nb

Substituting values of na and nb from (ii) & (iii)

⇒ 0.7 = 0.525P + 0.7P

⇒ 0.7 = 1.225 P

P= 0.7 / 1.225 =0.5714 atm

na = 0.525P = 0.525 * 0.5714 = 0.299 mol ≈ 0.3 mol

And,

nb = 0.7P = 0.7 * 0.5714 = 0.399 mol ≈ 0.4 mol